What is the simplified form of cos30degrees + (sin210degrees)(cos^2 150degrees)?
Answer by stanbon(75887)

What is the simplified form of cos30degrees + (sin210degrees)(cos^2 150degrees)?
-------------
cos(30) = [sqrt3]/2
sin(210) = -sin(30) = -1/2
cos(150)= -cos(30)= [-sqrt3]/2
---------------------
Your Problem:
{sqrt3]/2 + (-1/2)[[-sqrt3]/2]^2 = [sqrt3]/2 +(-1/2)(3/4)
= [sqrt3]/2 -(3/8)
= 0.8660-0.125
= 0.4910
============
Cheers,
Stan H.